what is the value of log 2

Grade:11

26 Answers

ganesh
10 years ago

the value of log2 is 3010

Nithin
10 years ago
log(2) is 0.30102
Indu
10 years ago
value of log2=0.3010. You can prove it.You could rewrite log(2) in various ways, such as: log(2) = log(10/5) = log(10) - log(5) = 1 - log(5) But that does not prove what it equals unless you have been given a value for log(5). You could use the definition of a logarithm to rewrite the expression: log(2) = 0.3010 <=> 10^(0.3010) = 2 And then use a calculator to verify that this produces an approximately true statement (since 0.3010 is a rounded number), but this does not qualify as a mathematical proof either.
RAHUL SINGH
10 years ago
log2 =0.3010
prince raj
10 years ago
0.3010
K RAJESH
10 years ago
log2 =0.30102
Adarsh
8 years ago
The value of log2=.3010
Vasanth Bhaskara
8 years ago
log2 = 0.30102
raghavi priya
8 years ago
The value of log2 is 0.3010
 
Nicho priyatham
8 years ago
0.3010
VAISHNAVI BATHINA
8 years ago
0
VAISHNAVI BATHINA
8 years ago
0.3010
SEELA YUVARAJU
8 years ago
0.301
KUNCHAM SAMPATH
8 years ago
Value of log 2 is 0.3010
sri tanish
8 years ago
the value of log2 is 0.301 . the literal meaning of this is if we multiply 10 with itself 0.301 times the answer is 2
Harshit Panwar
6 years ago
Value of log 2 is equal to 0.3010
Hope it helps you.
nivesh kesharwani
6 years ago
the value of log2 is 0.301
Ashna
6 years ago
log2=0.301
Ashna
6 years ago
log2=0.3010
Ashna
6 years ago
log2=0.3010
Ashna
6 years ago
log2=0.3
Ashna
6 years ago
Please give approval to any of the answer above .I think you have not even seen any of the answers the answer should be seen if you have asked it
Sai Soumya
5 years ago
The value of log 2 is 0.3010
You can also find the value of other numbers by getting a log table , which is easily accessible.
 
Hope this helps u
Sai Soumya
5 years ago
Dear Student ,
the value of log 2 is 0.301 .
You can also find the values of other numbers from the internet . Also , there is one log table available ,
 
Hope this helps you .
Have a good day .
Sayantan Garai
5 years ago
log2= 0.30102
Sayantan Garai
5 years ago
log2=0.30102999566398.